OpenStudy (anonymous):

HELP PLEASEEEEEEEE What are the solutions to the equation |2x -3| + 4 = 17? x = 8 only x = -8 or x = 5 x = -5 or x = 8 x = 12 only

OpenStudy (praetorian.10):

17 - 4 = 13

OpenStudy (praetorian.10):

- 3 = 12

OpenStudy (praetorian.10):

12 / 2 = 6

OpenStudy (praetorian.10):

therefore x must equal.................?

OpenStudy (campbell_st):

subtract 4 from both sides of the equation and you get \[\left| 2x - 3 \right| = 13\] now you have 2 equations to solve because it an absolute value 2x - 3 = 13 and 2x - 3 = -13 hope it helps

OpenStudy (praetorian.10):

or what that dude said

OpenStudy (praetorian.10):

i suppose YOLO

OpenStudy (campbell_st):

you can test you solutions by substitution.